OpenStudy (anonymous):

The average of three numbers is 69. The average of four different numbers is 90. What is the weighted mean of all the numbers together? 79 81 83 84

OpenStudy (kropot72):

The sum of the three numbers is 69 * 3 The sum of the four numbers is 90 * 4 The weighted mean is found by adding the above two results and dividing that total by 3 + 4 = 7\[weighted\ mean=\bar{x}=\frac{(69\times 3)+(90\times 4)}{7}=you\ can\ calculate\]

OpenStudy (anonymous):

81 :)

OpenStudy (kropot72):

Good work! Your answer is correct :)